OpenStudy (anonymous):

Solve for the variable. -4b = 15

OpenStudy (anonymous):

-60 60 -3.75 3.75

OpenStudy (anonymous):

-4b = 15 -4b / -4 = 15 / -4 b = -15 / 4 = -3.75

OpenStudy (uri):

b=15/-4 b=-15/4 =-3.75

OpenStudy (cggurumanjunath):